# Heegaard Floer Surgery Formula and Cosmetic Surgeries

### v1

@@ 3. Integer Surgeries
in particular, $\overline{\mathfrak{s}_{s}}$ is not one of the two special $\operatorname{Spin}^{c}$ structures and $-s>0$. By the above argument, $F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{+}=0$, hence also $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}=0$.
The conclusion follows by Theorem 2.3. ∎
3.1. Proof of Theorem for negative surgery coefficients
Continue to assume $n>0$, only we consider $Y_{-n}(K)$. In this case, let $W^{\prime}_{n(k-1)}(K)$ be the cobordism from $Y_{n(k-1)}(K)$ to $Y$. For each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/n{\mathbb{Z}}$,
$HF^{+}(Y_{-n}(K),i,\mathfrak{t})\cong H_{*}\left(MC\left(f_{2,i,\mathfrak{t}}^{+}:CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to\bigoplus^{nk}CF^{+}(Y,\mathfrak{t})\right)\right),$
@@ 3.1. Proof of Theorem for negative surgery coefficients
$f_{2,i,\mathfrak{t}}^{+}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k-1)}^{\prime}(K),\mathfrak{s}_{s}}^{+}:CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to CF^{+}(Y,\mathfrak{t}),$
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k-1)}(K))$ $|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k-1)\equiv 2s\mod 2n(k-1),$
$\text{and }s\equiv i\mod n\},$
$CF^{+}(Y_{n(k-1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k-1){\mathbb{Z}}\ |\ s\equiv i\mod n}CF^{+}(Y_{n(k-1)}(K),s,\mathfrak{t}).$
Consider the summand $CF^{+}(Y_{n(k-1)}(K),s,\mathfrak{t})$ for
$-\frac{n(k-1)}{2}\leq s<\frac{n(k-1)}{2},\hskip 14.22636pts\equiv i\mod n.$
Then, as above, if $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then $F_{W^{\prime}_{n(k-1)}(K),\mathfrak{s}_{s}}^{+}=0$. The conclusion again follows by Theorem 2.3.
4. Rational Surgeries
@@
$H_{*}(\mathbb{X}_{i,\mathfrak{t},p/q}^{+})\cong HF^{+}(Y_{p/q}(K),i,\mathfrak{t}).$
@@
By Kirby calculus, $Y_{p/q}(K)$ can be realized by surgery with coefficient $a$ inside the knot $K\#O_{q/r}\subset Y\#(-L(q,r))$, where $a$ is the greatest integer smaller than or equal to $p/q$, $a=\lfloor p/q\rfloor$, and $p/q=a+r/q$, and $O_{q/r}\subset-L(q,r)$ is the knot which is obtained by viewing one component of the the Hopf link as a knot inside the lens space $-L(q,r)$ thought of as $-q/r$ surgery on the other component of the Hopf link [OS10]. Thus, one obtains the rational surgery formula by proving a formula for Morse surgeries on rationally null-homologous knots and then applying a Kunneth principle for connected sums.
Analogously to the case of integer surgeries on null-homologous knots, for $K\subset Y$ a rationally null-homologous knot, $\lambda$ a framing on $K$, and $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, Ozsváth and Szabó define the chain map $D_{\mathfrak{s}}^{+}:{\mathbb{A}}_{\mathfrak{s}}^{+}(Y,\overline{K})\to{\mathbb{B}}_{\mathfrak{s}}^{+}(Y,\overline{K})$. Let ${\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$ be the mapping cone of $D_{\mathfrak{s}}^{+}$ [OS10].
Theorem 4.2 (Morse surgery formula [OS10]).
@@
$HF^{+}(Y_{\lambda}(K),\mathfrak{s})\cong H_{*}({\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$
@@
We recall the definition of $U$-knots from [OS10].
@@
A $U$-knot is a knot in a rational homology three-sphere $Y$ with the property that the induced filtration of $CF^{\infty}(Y,\mathfrak{s})$ is trivial. More precisely, given any $\xi\in\underline{\operatorname{Spin}^{c}}(Y,K)$, $CFK^{\infty}(Y,K,\xi)$ is chain homotopy equivalent via a relatively ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered homotopy to the chain complex $R$ which is a free, rank one ${\mathbb{Z}}[U,U^{-1}]$ module with the trivial differential.
Proposition 4.4.
@@
as ${\mathbb{Z}}\oplus{\mathbb{Z}}$-filtered chain complexes [OS10].
Lemma 4.5.
$O_{q/r}\subset-L(q,r)$ is a $U$-knot. There is an affine identification $\phi$ of the $\operatorname{Spin}^{c}$ structures on $\widehat{HFK}(L(q,r),O_{q,r})$ with ${\mathbb{Z}}$ such that
$\widehat{HFK}(L(q,r),O_{q,r},\phi(i))\cong\begin{cases}{\mathbb{Z}}&\text{if $0\leq i\leq q-1$,}\\ 0&\text{otherwise.}\end{cases}$
[OS10]
Let $K\subset Y$ be a null-homologous knot. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, let $H^{2}(Y,K,\mathfrak{t})\cong{\mathbb{Z}}$ denote the subgroup of $H^{2}(Y,K)$ corresponding to $\mathfrak{t}$, and also $H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})\cong{\mathbb{Z}}$.
Lemma 4.6.
@@
the following diagram commutes:
{tikzcd}
where $f(x,y)=qx+y$. Moreover, under this correspondence, if $K_{\lambda}$ is the pushoff of $K$ with respect to the integral framing $a$, then $\operatorname{PD}[K_{\lambda}]$ represents the element $p\in{\mathbb{Z}}\cong H^{2}(Y\#-L(q,r),K\#O_{q,r})$, where $p/q=a+r/q$.
@@ 5. Cosmetic Surgeries
The following lemma was proven in [Ni20].
@@
If $s\leq t$, then $*\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{t})_{*}$ and $*\operatorname{Im}(\widehat{h}_{s})_{*}\supset\operatorname{Im}(\widehat{h}_{t})_{*}$.
@@
There is an equivalence between $\widehat{v}_{s}$ and $\widehat{h}_{-s}$ for all $s$.
This equivalence is induced by changing the roles of the basepoints $w$ and $z$ in the doubly-pointed Heegaard diagram.
Lemma 5.3.
Let $g(K)$ be the genus of $K$, then for all $s\geq g(K)$, $\widehat{v}_{s}$, $\widehat{h}_{-s}$ induce isomorphisms on homology. For all $s\geq g(K)+1$, $\widehat{v}_{-s}$ and $\widehat{h}_{s}$ are nullhomotopic.
@@
Let $K\subset Y$ be a null-homologous knot. Suppose that $F$ is a taut Seifert surface for $K$. Let $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ for some $N$, and let $\iota:(Y,K)\to(X,K\times\{0\})$ be the inclusion map. Let $G\subset X$ be a properly-embedded smooth connected surface with $\partial G=K\times\{0\}$, $[G]=\iota_{*}[F]\in H^{2}(X,K\times\{0\})$, and $g(G)<g(F)$. Then
@@
Again, call a knot where such $(X,G)$ for $G$ a disk exists a blow-up unknotted knot. We also consider the unknot itself to be a blow-up unknotted knot.
@@
The proof makes use of the $CF^{\delta}$ variant of Heegaard Floer homology defined in [OS08]. For any nonnegative integer $\delta$, Denote by $CF^{\delta}(Y)$ the subcomplex of $CF^{+}(Y)$ which is the kernel of multiplication by $U^{\delta+1}$. Similarly, let $A_{s}^{\delta}$, $B^{\delta}$ denote the subcomplexes of $A_{s}^{+}$, $B^{+}$, respectively, which are the kernel of multiplication by $U^{\delta+1}$.
@@
(5.6) $0\to C\{i<0,j\geq g-1\}\to C\{i\geq 0\text{ or }j\geq g-1\}\xrightarrow{v_{g-1}^{+}}C\{i\geq 0\}\to 0$
The middle term is $A_{g-1}^{+}$, and $C\{i\geq 0\}=B^{+}$. Also,
$H_{*}(C\{i<0,j\geq g-1\})\cong H_{*}(C\{(-1,g-1)\}\cong H_{*}(C\{0,g\})=\widehat{HFK}(K,g)\neq 0.$
Therefore, $(v_{g-1}^{+})_{*}$ is not an isomorphism.
Suppose for the sake of contradiction that $(\widehat{v}_{g-1})_{*}$ is an isomorphism. Suppose also that $(v_{g-1}^{\delta})_{*}$ is an isomorphism for some $\delta\geq 0$. By [OS04a] Theorem 4.4, the diagram in Figure 1 commutes and the rows are exact, so by the five lemma, $(v_{g-1}^{\delta+1})_{*}$ is also an isomorphism. By induction, $(v_{g-1}^{\delta})_{*}$ is an isomorphism for all $\delta$. Thus, by Lemma 2.7 in [OS08], $(v_{g-1}^{+})_{*}$ is an isomorphism, which is a contradiction.
{tikzcd}
Figure 1.
@@
Let $K\subset Y$ be a blow-up unknotted knot, $p,q>0$ relatively prime, and let $x\in H_{*}(\widehat{A}_{j/q})$ for some $j$. If $\lfloor j/q\rfloor\geq 0$, there exists $\alpha\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ such that $(\widehat{D}_{p/q})_{*}(x+\alpha)=(\widehat{v}_{j/q})_{*}(x)$ (where we think of $H_{*}(\widehat{B}_{j/q})$ as including into $H_{*}(\widehat{\mathbb{B}})$). Similarly, if $\lfloor j/q\rfloor\leq 0$, there exists $\alpha\in H_{*}(\oplus_{i\leq j-p}\widehat{A}_{i/q})$ such that $(\widehat{D}_{p/q})_{*}(x+\alpha)=(\widehat{h}_{j/q})_{*}(x)$.
@@
For $s\geq 0$, we have $*\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s+1})_{*}\subset\dots$, and for $s\leq 0$, $*\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s-1})_{*}\subset\dots$. Thus, in the first case, there exists $y_{1}\in H_{*}(\widehat{A}_{(j+p)/q})$ such that $(\widehat{h}_{j/q})_{*}(x)=(\widehat{v}_{(j+p)/q})_{*}(y_{1})$. If $(\widehat{h}_{(j+p)/q})_{*}(y_{1})=0$, then stop and set $\alpha=y_{1}$, otherwise keep going. Inductively, given $y_{n}$ in $H_{*}(\widehat{A}_{(j+np)/q})$, there exists $y_{n+1}\in H_{*}(\widehat{A}_{(j+(n+1)p)/q})$ such that $(\widehat{h}_{(j+np)/q})_{*}(y_{n})=(\widehat{v}_{(j+(n+1)p)/q})_{*}$. By Lemma 5.3, this process eventually terminates and gives an element $\alpha=\sum_{n\geq 1}y_{n}\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ with the desired condition.
Similarly, in the case where $\lfloor j/q\rfloor\leq 0$, since $*\operatorname{Im}(\widehat{v}_{j/q})_{*}\subset\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, we can inductively add cancellation terms to the left. ∎
Using the mapping cone formula and the above lemmas, in the case where $K$ is blow-up unknotted, we may now calculate the rank of $\widehat{HF}(Y_{r}(K))$ for any positive rational number $r$.

### v2

@@ 3 Integer Surgeries
in particular, $\overline{\mathfrak{s}_{s}}$ is not one of the two special $\operatorname{Spin}^{c}$ structures and s>0-s>0. By the above argument, $F_{W_{n(k+1)}^{\prime}(K),\overline{\mathfrak{s}_{s}}}^{\delta}=0$, hence also $F_{W_{n(k+1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}=0$.
Therefore, $MC(f_{2,i,\mathfrak{t}}^{\delta})$ is quasi-isomorphic to the mapping cone of
$\sum_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}\left(f_{W_{n(k+1)}^{\prime}(K),\mathfrak{x}_{s}}^{\delta}+f_{W_{n(k+1)}^{\prime}(K),\mathfrak{y}_{s}}^{\delta}\right):CF^{\delta}(Y_{n(k+1)}(K),[i],\mathfrak{t})\to CF^{\delta}(Y,\mathfrak{t}).$
By Theorem 2.7, this chain complex is quasi-isomorphic to the mapping cone of
$\sum_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}):\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}A_{s}^{\delta}\to\bigoplus_{s\in{\mathbb{Z}}/n(k+1){\mathbb{Z}}\ :\ s\equiv i\pmod{n}}B^{\delta}.$
We rewrite this as the truncated mapping cone
$\sum_{-n(k+1)/2\leq s<n(k+1)/2\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}).$
We chose $k$ large enough such that $\frac{n(k+1)}{2}>g+\delta$, so by Lemma 2.6, the above mapping cone is quasi-isomorphic to the mapping cone of
$\sum_{s\ :\ s\equiv i\pmod{n}}(v_{s}^{\delta}+h_{s}^{\delta}):\bigoplus_{s\ :\ s\equiv i\pmod{n}}A_{s}^{\delta}\to\bigoplus_{s\ :\ s\equiv i\pmod{n}}B^{\delta},$
which by definition is ${\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$.
Therefore, we have the quasi-isomorphism
$CF^{\delta}(Y_{n}(K),i,\mathfrak{t})\cong{\mathbb{X}}_{i,\mathfrak{t}}^{\delta}(n)$
for all $\delta\geq 0$. Applying [OS08, Lemma 2.7], we conclude that
$CF^{+}(Y_{n}(K),i,\mathfrak{t})\cong{\mathbb{X}}_{i,\mathfrak{t}}^{+}(n).$
Proof of Theorem 3.3 for negative surgery coefficients.
Continue to assume $n>0$, only we consider $Y_{-n}(K)$. Again, we first work with the $\delta$ version for $\delta\geq 0$. In this case, let $W^{\prime}_{n(k-1)}(K)$ be the cobordism from $Y_{n(k-1)}(K)$ to $Y$. For each $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$, $i\in{\mathbb{Z}}/n{\mathbb{Z}}$,
$HF^{\delta}(Y_{-n}(K),i,\mathfrak{t})\cong H_{*}\left(MC\left(f_{2,i,\mathfrak{t}}^{\delta}:CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to\bigoplus^{nk}CF^{\delta}(Y,\mathfrak{t})\right)\right),$
@@
$f_{2,i,\mathfrak{t}}^{\delta}=\sum_{\mathfrak{s}_{s}\in\mathcal{X}_{i}}f_{W_{n(k-1)}^{\prime}(K),\mathfrak{s}_{s}}^{\delta}:CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})\to CF^{\delta}(Y,\mathfrak{t}),$
$\mathcal{X}_{i}=\{\mathfrak{s}_{s}\in\operatorname{Spin}^{c}(W^{\prime}_{n(k-1)}(K))\$$|\ \mathfrak{s}_{s}|_{Y}=\mathfrak{t},$
$\langle c_{1}(\mathfrak{s}_{s}),[\widehat{F}]\rangle-n(k-1)\equiv 2s\pmod{2n(k-1)}$,
$\text{and }s\equiv i\pmod{n}\}$,
$CF^{\delta}(Y_{n(k-1)}(K),[i],\mathfrak{t})=\bigoplus_{s\in{\mathbb{Z}}/n(k-1){\mathbb{Z}}\ |\ s\equiv i\pmod{n}}CF^{\delta}(Y_{n(k-1)}(K),s,\mathfrak{t}).$
Consider the summand $CF^{\delta}(Y_{n(k-1)}(K),s,\mathfrak{t})$ for
$\frac{n(k-1)}{2}\leq s<\frac{n(k-1)}{2},\hskip 14.22636pts\equiv i\pmod{n}$.
Then, as above, if $\mathfrak{s}_{s}\in\mathcal{X}_{i}\setminus\{\mathfrak{x}_{s},\mathfrak{y}_{s}\}$, then $F_{W^{\prime}_{n(k-1)}(K),\mathfrak{s}_{s}}^{\delta}=0$. The conclusion follows again by the same argument as in the positive surgery case.
4 Rational Surgeries
@@
$H_{*}(\mathbb{X}_{p/q,i,\mathfrak{t}}^{+})\cong HF^{+}(Y_{p/q}(K),i,\mathfrak{t}).$
@@
Let $a=\lfloor p/q\rfloor$ and $p/q=a+r/q$. Consider the Hopf link in $S^{3}$ and perform q/r-q/r surgery on one component of the link. Let $O_{q/r}\subset-L(q,r)$ be the other component of the link viewed as a knot in the resulting lens space after surgery. By Kirby calculus, $Y_{p/q}(K)$ can be realized by surgery with coefficient $a$ inside the knot $K\#O_{q/r}\subset Y\#(-L(q,r))$ [OS10]. Thus, one obtains the rational surgery formula by proving a formula for Morse surgeries on rationally null-homologous knots and then applying a Kunneth principle for connected sums.
Analogously to the case of integer surgeries on null-homologous knots, for $K\subset Y$ a rationally null-homologous knot with an orientation (we use $\overline{K}$ to denote the oriented knot), $\lambda$ a framing on $K$, and $\mathfrak{s}\in\operatorname{Spin}^{c}(Y_{\lambda}(K))$, define the chain map $D_{\mathfrak{s}}^{+}:{\mathbb{A}}_{\mathfrak{s}}^{+}(Y,\overline{K})\to{\mathbb{B}}_{\mathfrak{s}}^{+}(Y,\overline{K})$ as in [OS10] Theorem 6.1. Let ${\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda)$ be the mapping cone of $D_{\mathfrak{s}}^{+}$ [OS10].
Theorem 4.2 (Morse surgery formula [OS10] Theorem 6.1).
@@
$HF^{+}(Y_{\lambda}(K),\mathfrak{s})\cong H_{*}({\mathbb{X}}_{\mathfrak{s}}^{+}(Y,\overline{K},\lambda))$
@@
We recall the definition of $U$-knots from [OS10] Definition 5.2.
@@
[…]
Proposition 4.4 ([OS10], Corollary 5.3).
@@
[…]
Lemma 4.5 ([OS10], Lemma 7.1).
$O_{q/r}\subset-L(q,r)$ is a $U$-knot. There is an affine identification $\phi$ of the $\operatorname{Spin}^{c}$ structures on $\widehat{HFK}(L(q,r),O_{q/r})$ with $\mathbb{Z}$} such that
$\widehat{HFK}(L(q,r),O_{q/r},\phi(i))\cong\begin{cases}{\mathbb{Z}}&\text{if $0\leq i\leq q-1$,}\\ 0&\text{otherwise.}\end{cases}$
Let $K\subset Y$ be a null-homologous knot. As in section 2.2 of [OS10], $H^{2}(Y,K)$ is defined to be $H^{2}(Y-\operatorname{nd}(K),\partial\operatorname{nd}(K))$ and is an affine space for $\underline{\operatorname{Spin}^{c}}(Y,K)$. There is a natural map $\underline{\operatorname{Spin}^{c}}(Y,K)\to\operatorname{Spin}^{c}(Y)$ which allows us to identify
$\operatorname{Spin}^{c}(Y)\cong\frac{\underline{\operatorname{Spin}^{c}}(Y,K)}{{\mathbb{Z}}\cdot\operatorname{PD}[\mu]}.$
Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ over $Y$, and under the above identification, let $H^{2}(Y,K,\mathfrak{t})\cong{\mathbb{Z}}$ denote the subgroup of $H^{2}(Y,K)$ corresponding to $\mathfrak{t}$, and also $H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})\cong{\mathbb{Z}}$.
With this notation, the following more general version of [OS10, Lemma 7.2] holds with analogous proof.
Lemma 4.6 ([OS10], Lemma 7.2).
@@
the diagram in Figure 2 commutes where $f(x,y)=qx+y$. Moreover, under this correspondence, if $K_{\lambda}$ is the pushoff of $K$ with respect to the integral framing $a$, then $\operatorname{PD}[K_{\lambda}]$ represents the element $p\in{\mathbb{Z}}\cong H^{2}(Y\#-L(q,r),K\#O_{q,r})$, where $p/q=a+r/q$.
$\mathbb{Z}\oplus\mathbb{Z}$$\mathbb{Z}$$H^{2}(Y,K,\mathfrak{t})\oplus H^{2}(-L(q,r),O_{q,r})$$H^{2}(Y\#-L(q,r),K\#O_{q,r},\mathfrak{t})$$\scriptstyle{f}\scriptstyle{\cong}\scriptstyle{\cong}$
Figure 2:
@@ 5 Cosmetic Surgeries
The following lemma was proven in [NI20] Equations (5) and (6).
@@
If $s\leq t$, then $\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{t})_{*}$ and $\operatorname{Im}(\widehat{h}_{s})_{*}\supset\operatorname{Im}(\widehat{h}_{t})_{*}$.
As a consequence of the hat version of Lemma 2.5, $\operatorname{rank}H_{*}(\widehat{A}_{s})=\operatorname{rank}H_{*}(\widehat{A}_{-s})$ and $\operatorname{rank}(\widehat{v}_{s})_{*}=\operatorname{rank}(\widehat{h}_{-s})_{*}$ for all $s$.
@@
Suppose for some $s\in{\mathbb{Z}}$ that $\widehat{v}_{s}$ or $\widehat{h}_{s}$ induces an isomorphism on homology. Then $v_{s}^{+}$ or $h_{s}^{+}$ also induces an isomorphism on homology, respectively.
Proof.
We prove this lemma for $\widehat{v}_{s}$, and the case for $\widehat{h}_{s}$ is proven similarly. We will show by induction that $(v_{s}^{\delta})_{*}$ is an isomorphism for all $\delta\geq 0$, which by Lemma 2.7 in [OS08] implies that $(v_{g-1}^{+})_{*}$ is an isomorphism. Suppose that $(v_{s}^{\delta})_{*}$ is an isomorphism for some $\delta\geq 0$ (the base case $\delta=0$ is $v_{s}^{0}=\widehat{v}_{s}$).
Clearly, the maps $\widehat{v}_{s}$, $v_{s}^{\delta+1}$, and $v_{s}^{\delta}$ fit into the commutative diagram of short exact sequences in Figure 3. Therefore, the diagram in Figure 4 commutes and the rows are exact, so by the five lemma, $(v_{s}^{\delta+1})_{*}$ is also an isomorphism. ∎
[…]
Figure 3:
$H_{*}(A_{s}^{\delta})$$H_{*}(\widehat{A}_{s})$$H_{*}(A_{s}^{\delta+1})$$H_{*}(A_{s}^{\delta})$$H_{*}(\widehat{A}_{s})$$H_{*}(B^{\delta})$${H_{*}(\widehat{B})}$$H_{*}(B^{\delta+1})$$H_{*}(B^{\delta})$${H_{*}(\widehat{B})}\scriptstyle{\partial}$$\scriptstyle{(v_{s}^{\delta})_{*}}$$\scriptstyle{i_{*}}$$\scriptstyle{(\widehat{v}_{s})_{*}}$$\scriptstyle{U_{*}}$$\scriptstyle{(v_{s}^{\delta+1})_{*}}\scriptstyle{\partial}$$\scriptstyle{(v_{s}^{\delta})_{*}}$$\scriptstyle{(\widehat{v}_{s})_{*}}\scriptstyle{\partial}$$\scriptstyle{i_{*}}$$\scriptstyle{U_{*}}\scriptstyle{\partial}$
Figure 4:
@@
Let $K\subset Y$ be a null-homologous knot. Suppose that $F$ is a taut Seifert surface for $K$. Let $X=(Y\times[0,1])\#N\overline{{\mathbb{C}}P^{2}}$ for some $N$, and let $\iota:(Y,K)\to(X,K\times\{0\})$ be the inclusion map. Let $G\subset X$ be a properly-embedded smooth connected surface with $\partial G=K\times\{0\}$, $G]=\iota_{*}[F]\in H^{2}(X,K\times\{0\})$, and $g(G)<g(F)$. Then
@@
The same is true for the hat version of knot Floer homology by essentially the same proof.
Again, call a knot where such $(X,G)$ for $G$ a disk exists a blow-up bounding knot. We also consider the unknot itself to be a blow-up bounding knot.
The following argument is due to Ni:
Lemma 5.4.
[…]
Proof.
Denote by $O$ the unknot in $Y$. There is a sequence of $N$ crossing changes changing a positive crossing to a negative one that changes $O$ to the knot $K$. Let $W_{n}^{\prime}(K):Y_{n}(K)\to Y$ and $W_{n}^{\prime}(O):Y_{n}(O)\to Y$ denote the natural two-handle cobordisms. Fix a $\operatorname{Spin}^{c}$ structure $\mathfrak{t}$ on $Y$, let $\mathfrak{x}_{0}(O)\in\operatorname{Spin}^{c}(W_{n}^{\prime}(O))$ and $\mathfrak{x}_{0}(K)\in\operatorname{Spin}^{c}(W_{n}^{\prime}(K))$ as in Section 2 restrict to the $\operatorname{Spin}^{c}$ strctures $\mathfrak{x}_{0}(O)|Y=\mathfrak{x}_{0}(K)|Y=\mathfrak{t}$ and $\mathfrak{t}_{0}(O)=\mathfrak{x}_{0}(O)|Y_{n}(O)$, $\mathfrak{t}_{0}(K)=\mathfrak{x}_{0}(K)|Y_{n}(K)$. Let $\mathfrak{S}$ be the $\operatorname{Spin}^{c}$ structure on $W_{n}^{\prime}(O)\#N\overline{{\mathbb{C}}P^{2}}$ given by the connected sum of $\mathfrak{x}_{0}(O)$ with $N$ copies of $\mathfrak{s}_{1}\in\operatorname{Spin}^{c}(\overline{{\mathbb{C}}P^{2}})$, where $\mathfrak{s}_{1}$ satisfies that $c_{1}(\mathfrak{s}_{1})$ generates $H^{2}(\overline{{\mathbb{C}}P^{2}})$.
By Kirby calculus, the Kirby diagram of the $n$-framed unknot in $Y$ blown up $N$ times is equivalent to the diagram of the $n$-framed knot $K$ in $Y$. Thus, the cobordism $W_{n}^{\prime}(O)\#N\overline{{\mathbb{C}}P^{2}}$ is the composition of two cobordisms
$Y_{n}(O)\xrightarrow{X_{n}(K)}Y_{n}(K)\xrightarrow{W_{n}^{\prime}(K)}Y$
for some cobordism $X_{n}(K)$. The $\operatorname{Spin}^{c}$ structure $\mathfrak{S}$ restricts to $\mathfrak{t}_{0}(O)$ on $Y_{n}(O)$, $\mathfrak{t}_{0}(K)$ on $Y_{n}(K)$, and $\mathfrak{t}$ on $Y$. Therefore, $X_{n}(K)$ induces a cobordism map
$\Phi=\widehat{F}_{X_{n}(K),\mathfrak{S}|X_{n}(K)}:\widehat{HF}(Y_{n}(O),\mathfrak{t}_{0}(O))\to\widehat{HF}(Y_{n}(K),\mathfrak{t}_{0}(K))$
that by Theorem 2.7 fits into the commutative diagram in Figure 5. By a sample calculation, $(\widehat{v}_{0}(O))_{*}$ is surjective. Therefore $(\widehat{v}_{0}(K))_{*}$ must also be surjective. Direct summing over all $\mathfrak{t}\in\operatorname{Spin}^{c}(Y)$ proves the statement.
$\widehat{HF}(Y_{n}(O),\mathfrak{t}_{0}(O))$$\widehat{HF}(Y_{n}(K),\mathfrak{t}_{0}(K))$$H_{*}(\widehat{A}_{0,\mathfrak{t}}(O))$$H_{*}(\widehat{A}_{0,\mathfrak{t}}(K))$$H_{*}(\widehat{B}_{\mathfrak{t}}(O))$$H_{*}(\widehat{B}_{\mathfrak{t}}(K))$$\scriptstyle{\Phi}\scriptstyle{\cong}\scriptstyle{\cong}$$\scriptstyle{(\widehat{v}_{0}(O))_{*}}$$\scriptstyle{(\widehat{v}_{0}(K))_{*}}$$\scriptstyle{=}$
Figure 5:
@@
$0\to C\{i<0,j\geq g-1\}\to C\{i\geq 0\text{ or }j\geq g-1\}\xrightarrow{v_{g-1}^{+}}C\{i\geq 0\}\to 0.$ (5.6)
The middle term is $A_{g-1}^{+}$, and $C\{i\geq 0\}=B^{+}$. The assumption that $C$ is reduced together with the $U$-isomorphism and the fact that knot Floer homology detects genus implies that $C\{i,j\}$ is nonzero only if $g\leq i-j\leq g$. Therefore,
$H_{*}(C\{i<0,j\geq g-1\})\cong H_{*}(C\{(-1,g-1)\}),$
which by the $U$-isomorphism is isomorphic to
$H_{*}(C\{0,g\})=\widehat{HFK}(K,g),$
and $\widehat{HFK}(K,g)\neq 0$ since knot Floer homology detects genus. Therefore, $(v_{g-1}^{+})_{*}$ is not an isomorphism. By Lemma 5.2, $(\widehat{v}_{g-1})_{*}$ is also not an isomorphism.
We abuse notation to let $\widehat{A}_{j/q}$ denote $\widehat{A}_{\lfloor j/q\rfloor}$, and also $\widehat{B}_{j/q}$, $\widehat{v}_{j/q}$, $\widehat{h}_{j/q}$ similarly.
@@
[…]
@@
Since $K$ is blow-up bounding, by Lemma 5.3, for $s\geq 0$, we have $\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{v}_{s+1})_{*}\subset\dots$, and for $s\leq 0$, $\operatorname{Im}(\widehat{v}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s})_{*}\subset\operatorname{Im}(\widehat{h}_{s-1})_{*}\subset\dots$. Thus, in the first case, there exists $y_{1}\in H_{*}(\widehat{A}_{(j+p)/q})$ such that $(\widehat{h}_{j/q})_{*}(x)=(\widehat{v}_{(j+p)/q})_{*}(y_{1})$. If $(\widehat{h}_{(j+p)/q})_{*}(y_{1})=0$, then stop and set $\alpha=y_{1}$, otherwise keep going. Inductively, given $y_{n}$ in $H_{*}(\widehat{A}_{(j+np)/q})$, there exists $y_{n+1}\in H_{*}(\widehat{A}_{(j+(n+1)p)/q})$ such that $(\widehat{h}_{(j+np)/q})_{*}(y_{n})=(\widehat{v}_{(j+(n+1)p)/q})_{*}$. By Lemma 2.6, this process eventually terminates and gives an element $\alpha=\sum_{n\geq 1}y_{n}\in H_{*}(\oplus_{i\geq j+p}\widehat{A}_{i/q})$ with the desired condition.
Similarly, in the case where $\lfloor j/q\rfloor\leq 0$, since $\operatorname{Im}(\widehat{v}_{j/q})_{*}\subset\operatorname{Im}(\widehat{h}_{(j-p)/q})_{*}$, we can inductively add cancellation terms to the left. ∎
Using the mapping cone formula and the above lemmas, in the case where $K$ is blow-up bounding, we may now calculate the rank of $\widehat{HF}(Y_{r}(K))$ for any positive rational number $r$.